If McDonald’s spent $15,000 of the total Quick Serve Restaurant industry media spend of $100,000, what % is McDonald’s media spend?

a.

5%

b.

10%

c.

15%

d.

20%

e.

25%
Answer #1

Answer: c. 15%

McDonald's media spent = $15000

Total quick serve restaurant industry's media spent = $100000

% of McDonald's media spent = ($15000/$100000)*100% = 0.15*100% = 15%